Rank the following pairs of force according to the magnitude of the net force, with the largest magnitude net force at the top and the smallest at the bottom. t Place these in the proper order (9 N, left) and (6 N, right) (5 N, right) and (3 N, right) (9 N, right) and (9 N, left) (7 N, right) and (8 N, left)
Answer #1

A) 9N - 6N = 3N (left)

B) 5N + 3N = 8N (right)

C) 9N - 9N = 0

D) 7N - 8N = 1N (left)

B > A > D > C

Comment in case any doubt please rate my answer ....